OpenStudy (anonymous):

Change the given equation into slope-intercept form. 2x – 5y = -10

OpenStudy (anonymous):

slope intercept form is y = mx + b. 2x - 5y = -10 (subtract 2x from both sides) -5y = -2x - 10 (multiply the equation by -1 to make y positive) 5y = 2x + 10 (now divide both sides by 5) y = 2/5x + 2 any questions ?

OpenStudy (anonymous):

so -5/2

OpenStudy (anonymous):

-2/5

OpenStudy (anonymous):

slope intercept form is y = mx + b your answer is y = 2/5x + 2